# Liming Liu1    Zhenghao Xu1    Zixuan Zhang1    Hao Kang1 Zichong Li1    Chen Liang2    Weizhu Chen2    Tuo Zhao1

### v2

@@ A.2 Ablation of $r$ and $\gamma$ on 130M model
For the ablation experiments on COSMOS in Section 4.1, we tune discount factor $\gamma$ and rank $r$ together to show COSMOS isn’t very sensitive to hyperparameters. We fix the learning rate for embedding and output layers to be $32e-$ 3, fix the learning rate for COSMOS to be 5e-4, and sweep over the cross product of $r\in\{32,64,128\}$ and $\gamma\in\{0.1,0.25,0.5,1\}$. With all these hyperparameters COSMOS outputs comparable results to MUON.
@@ A.3 Ablation of normalization
[…]
@@ A.4 Profiling experiments
Appendix B Two-sided COSMOS
Similar to two-sided SOAP in Vyas et al. (2024), we provide a two-sided variant of COSMOS in Algorithm 4.
@@ Appendix B Two-sided COSMOS
Appendix C Additional Experiments
@@
C.1 Additional Experiments for Normalization Ablation
We conduct additional experiments on LLaMA-350M to further validate that normalization is not the main reason for COSMOS outperforming MUON. As shown in Figure 6, normalization does not make much difference to MUON, while COSMOS consistently outperforms both of them.
Figure 6: Comparison of COSMOS, MUON and MUON with normalization for LLaMA-350M on C4. MUON’s performance does not differ by much with normalization, implying that normalization is not the dominant reason why COSMOS outperforms MUON.
C.2 Additional Experiments on WikiText
This section provides additional experiments on WikiText (Merity et al., 2016) and GPT-2 (Radford et al., 2019). To be more specific, we train GPT2-small(125M) and GPT2-medium (355M) on the Wikitext-103 dataset. We discard learnable position embeddings and use RoPE (Su et al., 2024) as a replacement. For GPT2-small, we use learning rate 4e-3 for the embedding layer and 4e-4 for MUON/COSMOS. For GPT2-medium, we use learning rate 2e-3 for the embedding layer and 5e-4 for MUON/COSMOS. $\gamma$ is still set to be the ratio of the hidden layer learning rate to the embedding layer learning rate.
@@ C.2 Additional Experiments on WikiText
The results for GPT2-small and GPT2-medium are provided in Figures 7 and 8, respectively. We observe that COSMOS consistently outperforms MUON, showing that it does not overfit any particular model or dataset.
Figure 7: GPT2-small trained on WikiText-103 dataset. COSMOS consistently outperforms MUON.
Figure 8: GPT2-medium trained on WikiText-103 dataset. COSMOS consistently outperforms MUON.
C.3 Smaller learning rate for MUON/COSMOS on LLaMA-1B
[…]

### v3

@@ A.2 Ablation of $r$ and $\gamma$ on 130M model
For the ablation experiments on COSMOS in Section 4.1, we tune discount factor $\gamma$ and rank $r$ together to show COSMOS isn’t very sensitive to hyperparameters. We fix the learning rate for embedding and output layers to be 2e32e-3, fix the learning rate for COSMOS to be 5e-4, and sweep over the cross product of $r\in\{32,64,128\}$ and $\gamma\in\{0.1,0.25,0.5,1\}$. With all these hyperparameters COSMOS outputs comparable results to MUON.
@@ A.3 Ablation of normalization
As mentioned in the normalization paragraph in Section 4.1, to exclude the possibility that the better performance of COSMOS than MUON is simply because the normalization function NORM, we modify the normalization method of MUON to be NORM and rerun the experiments for 130M and 350M models. We still tune the learning rate of MUON ++ NORM on $\{$5e-3, 1e-2, 2e-2, 4e-2}\}, and present their best performance.
@@ A.4 Profiling experiments
Appendix B Experiments Details on Modded-NanoGPT
As discussed in Section 4.3, we directly use Muon’s reproducible logs on modded NanoGPT. In the setting of GPT-2 Small (124M), they set the learning rate for embedding layer (optimized with Adam) to be 3.6e-3, and the learning rate for hidden layers (optimized with Muon) to be 3.6e-4. Also they use Warmup-Stable-Decay (WSD) schedule instead of Cosine schedule. Their batch size is 512, sequence length is 1024 and number of iterations is 6200.
In the training of 124M model, we followed their original setting for Muon and only additionally searched $\beta_{1}$ within $\{0.9,0.95\}$. For COSMOS, we adopted the same setting as Muon, also searching $\beta_{1}$ in $\{0.9,0.95\}$. For $\beta_{2}$ and $\gamma$ in COSMOS, we set them to $0.95$ and $0.2$ without additional search.
In the training of GPT-2 Medium (350M), they used a very uncommon setting: the learning rate for embedding layer is 0.3, which is very large. But the learning rate for output layer is still 3e-3. They also reduced the momentum for the embedding layer and output layer to $0.8$ – which is also an uncommon choice.
To demonstrate the generality of COSMOS across various settings, our experiments on the 350M model completely follow their setting. We only additionally search $\beta_{1}$ within $\{0.9,0.95\}$ for both Muon and COSMOS. However, since this setting is indeed uncommon and subsequent work on Muon scaling (Shah et al., 2025; Liu et al., 2025) did not follow it, we did not adopt this setting in our other experiments (LLaMA on C4 and GPT2 on Wikitext103).
Appendix C Two-sided COSMOS
For simplicity, we only consider the one-side version of COSMOS in this paper. Like SOAP, COSMOS can be further generalized to a two-sided version. Similar to two-sided SOAP in Vyas et al. (2024), we provide a two-sided variant of COSMOS in Algorithm 4.
@@ Appendix C Two-sided COSMOS
Appendix D Additional Experiments
@@
D.1 Experiments for Normalization Ablation
We conduct additional experiments on LLaMA-130M and LLaMA-350M to validate that normalization is not the main reason for COSMOS outperforming MUON. As shown in Figure 5, normalization does not make much difference to MUON, while COSMOS consistently outperforms both of them.
(a) Comparison of COSMOS, MUON and MUON with normalization for LLaMA-130M on C4.
(b) Comparison of COSMOS, MUON and MUON with normalization for LLaMA-350M on C4.
Figure 5: Comparison of COSMOS, MUON, and MUON with normalization on LLaMA-130M and LLaMA-350M for C4.
D.2 Experiments on WikiText
This section discuss the details of the experiments on WikiText (Merity et al., 2016) and GPT-2 (Radford et al., 2019). To be more specific, we train GPT2-small(125M) and GPT2-medium (355M) on the Wikitext-103 dataset. We discard learnable position embeddings and use RoPE (Su et al., 2024) as a replacement.
For GPT2-small, we tune the learning rate of Adam on $\{$5e-4, 1e-3, 2e-3, 4e-3, 8e-3}\}, and find 4e-3 is the optimal learning rate for Adam. Then in MUON/COSMOS, we use learning rate 4e-3 for the embedding layer and 4e-4 for MUON/COSMOS.
Similarly, for GPT2-medium, we tune the learning rate of Adam on $\{$5e-4, 1e-3, 2e-3, 4e-3, 8e-3}\}, and find 2e-3 is the optimal learning rate for Adam. Then in MUON/COSMOS, we use learning rate 2e-3 for the embedding layer and 5e-4 for MUON/COSMOS.
For COSMOS, $\gamma$ is still set to be the ratio of the hidden layer learning rate to the embedding layer learning rate in both models.
@@ D.2 Experiments on WikiText
The results for GPT2-small and GPT2-medium are provided in Figures 4(a) and 4(b), respectively. We observe that COSMOS consistently outperforms MUON, showing that it does not overfit any particular model or dataset.
D.3 Smaller learning rate for MUON/COSMOS on LLaMA-1B
As discussed in section A.1.3, we tune the learning rate for MUON/COSMOS on $\{$2e-4, 5e-4}\}. We find 5e-4 is better and present its corresponding results in the main text. Here we present the result for 2e-4 in Figure 6, where COSMOS still outperforms MUON.
@@ D.3 Smaller learning rate for MUON/COSMOS on LLaMA-1B
D.4 Wall-Clock time plot for LLaMA-1B
Based on the throughput we calculate in Table 6, we rescale the X-axis of Figure 2(b) to be wall-clock time and present the result in Figure 7.
Figure 7: Wall-Clock time plot for our training on LLaMA-1B.
D.5 Experiments on shorter sequences
To validate the correctness of our implementation, we compare GaLore with COSMOS and Adam on 256 sequence length as adopted in Zhao et al. (2024a). As shown in Figure 8, GaLore and Adam are more comparable in shorter sequence setting, suggesting our implementation is correct and the degradation of GaLore shown in Figure 1 and Table 2 is mainly due to long sequence length.